You need to buy 8 packages of meat. A package of hotdogs costs $1.89 and package
of hamburgers costs $5.19. You spend a total of $31.62.
a) Let x represent the number of packages of hotdogs bought and let y represent
the number of packages of hamburgers bought. Write a system of equations you
could solve to find the number of packages of each type of meat bought.
b) solve the system

Answer by Paul(988)

189x+519(8-x)=3162
189x+4152-519x=3162
-330x=-990
x=3
8-3 = 5
Hence, for hotdogs you need 3 packages of meat and for hamburgers you need 5 packages of meat